# DESI forecast for Dark Matter-Neutrino interactions using EFTofLSS

### v1

@@ 1 Introduction
So far, the standard $\Lambda$CDM model of cosmology has proved to be a very good description of our universe [1], with some caveats such as e.g., the $H_{0}$ and $\sigma_{8}$ tensions [2] and the recent dark energy hints from DESI [3]. However, this model relies on two components whose natures are not well understood: dark energy, which in the standard picture takes the form of a cosmological constant $\Lambda$, and dark matter (DM), which in the baseline model is assumed to be non-interacting and cold. Of course, the assumption of the cold dark matter (CDM) is not truly a physical description of the DM particle, but rather a description of its apparent macroscopic behaviour. A significant amount of effort has been devoted to proposing particle physics models that can satisfy both cosmological observations and laboratory constraints, while also being produced in suitable quantities in the early universe. [4, 5, 6, 7, 8]. Such models will often lead to non-zero couplings with standard model particles, which could potentially give rise to signals in our experiments, or alternatively constrain the possible parameter space through the absence of detections [9, 10, 11, 12, 13, 14]. Cosmological observations have also been key in constraining the possible parameter space of different types of dark matter models, setting limits on interactions with baryons [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], photons [15, 27, 17, 28, 29, 20, 19, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 26], neutrinos [15, 17, 40, 41, 42, 34, 43, 44, 45, 46, 47], dark radiation species [48, 49, 50, 51, 52, 53, 54, 55, 56, 26, 57], DM self-interactions [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 15, 17], as well as warm DM models [68, 69, 70, 71, 72, 73, 74, 75, 76, 77], and axion DM models [78, 79, 80, 81, 82].
@@ 2 Dark matter-neutrino interactions
In this work we assume a scattering interaction between dark matter and standard model neutrinos, as described in Refs. [42, 43, 44]. A coupling between neutrinos and dark matter has been suggested as a possible source of the non-zero neutrino masses, and such couplings are exceedingly difficult to probe in laboratory experiments because of the weak coupling between neutrinos and other standard model particles [114, 115, 116]. We use the approximation of massless neutrinos for the sake of simplicity in the EFTofLSS calculations, but this is unlikely to have a large impact on the final constraints, as the inclusion of neutrino masses has been shown to only have minor effects on the bounds of the interaction strength [45, 47]. We restrict out analysis to the simplest form of interaction, using a temperature independent cross section.
@@
$u_{\nu\chi}\equiv\frac{\sigma_{0}}{\sigma_{\mathrm{Th}}}\left(\frac{m_{\chi}}{100\,\mathrm{GeV}}\right)^{-1},$ (2.1)
where $\sigma_{\mathrm{Th}}\approx 6.65\times 10^{-29}\,\mathrm{m}^{2}$ is the Thomson scattering cross-section, and $m_{\chi}$ is the dark matter particle mass. The mass factor is included in the definition of the interaction strength to avoid a parameter degeneracy between the scattering cross section and the DM particle mass.
Figure 1: Linear and EFTofLSS matter power spectrum for two IDM scenarios, with 50% and 90% IDM, respectively, as well as standard CDM for reference. For the nonlinear power spectrum we have set counterterms to zero. It is evident that the scenario with the larger IDM fraction is more suppressed, but for both models the suppression starts at the same scale, as they have the same interaction strength $u_{\nu\chi}$.
@@ 3 Effective Field Theory of Large Scale Structure
For the definition of EFTofLSS parameters, we rely on the CLASS-OneLoop parameterization, which is based on the convention in [120] and detailed in [108]. When modelling the one-loop galaxy power spectrum in real space at a given redshift, we have four bias parameters $(b_{1},b_{2},b_{{\cal G}_{2}},b_{\Gamma_{3}})$, one counter-term $c_{0}$, and two stochastic parameters $(s_{0},s_{1})$ For the one-loop galaxy power spectrum wedges in redshift space, we have the same four bias parameters, four counter-terms $(c_{0},c_{1},c_{2},c_{3})$, and four stochastic terms $(s_{0},s_{1},s_{2},s_{3})$.333[108] showed that $s_{1}$ is mostly degenerate with other parameters. Thus, in the following, we can limit ourselves to varying the three parameters $(s_{0},s_{2},s_{3})$. Note that $s_{3}$ is often refereed to as $s_{4}$ in other works.
@@
with flat uninformative priors on $(\tilde{b}_{1},\tilde{b}_{2},\tilde{b}_{{\cal G}2},\tilde{b}_{\Gamma_{3}})$. The fiducial values of $b_{{\cal G}_{2}},b_{\Gamma_{3}}$ are chosen according to the co-evolution model while $b_{2}$ follows a polynomial fit; this is the same approach as used in [108]. For the seven counter-terms and stochastic parameters, we compare three assumptions:
@@ 4.2 Comparison with intermediate and pessimistic cases
Overall, the impact of loosening our assumptions on stochastic terms and counterterms indicates that there is a significant degeneracy between the interaction parameter $u_{\nu\chi}$ and the stochastic contributions or counterterms. Our results in the intermediate and the pessimistic cases indicate that in the absence of well-motivated informative priors on these nuisance parameters, the limits on dark matter-neutrino interactions from galaxy clustering power spectrum using the EFTofLSS model will likely not to be competitive with those from the CMB. This conclusion is expected to be extendable to other models with small-scale suppression of linear matter power spectrum. However, the constraining power observed in the optimistic case suggests that improved constraints may be achievable in the future provided that a physically-motivated model for the redshift dependence of the EFTofLSS parameters can be established.
@@ Acknowledgments
The authors would like to thank Felix Kahlhöfer for useful discussions. The authors gratefully acknowledge the computing time provided to them on the high-performance computer Lichtenberg at the NHR Center NHR4CES@TUDa, funded by the German Federal Ministry of Education and Research (BMBF) and the Hessian Ministry of Science and Research, Art and Culture (HMWK). JL and MM acknowledge support from the DFG grant LE 3742/8-1. Some tests of the likelihood were performed with computing resources granted by RWTH Aachen University under projects thes1654 and rwth1389. SC acknowledges support by the Science and Technology Facilities Council (grant number ST/W001225/1).

### v2

@@ 1 Introduction
So far, the standard $\Lambda$CDM model of cosmology has proved to be a very good description of our universe [1], with some caveats such as e.g., the $H_{0}$ and $\sigma_{8}$ tensions [2] and the recent dark energy hints from DESI [3]. However, this model relies on two components whose natures are not well understood: dark energy, which in the standard picture takes the form of a cosmological constant $\Lambda$, and dark matter (DM), which in the baseline model is assumed to be non-interacting and cold. Of course, the assumption of the cold dark matter (CDM) is not truly a physical description of the DM particle, but rather a description of its apparent macroscopic behaviour. A significant amount of effort has been devoted to proposing particle physics models that can satisfy both cosmological observations and laboratory constraints, while also being produced in suitable quantities in the early universe. [4, 5, 6, 7, 8]. Such models will often lead to non-zero couplings with standard model particles, which could potentially give rise to signals in our experiments, or alternatively constrain the possible parameter space through the absence of detections [9, 10, 11, 12, 13, 14]. Cosmological observations have also been key in constraining the possible parameter space of different types of dark matter models, setting limits on interactions with baryons [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], photons [15, 27, 17, 28, 29, 20, 19, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 26], neutrinos [15, 17, 40, 41, 42, 34, 43, 44, 45, 46, 47, 48], dark radiation species [49, 50, 51, 52, 53, 54, 55, 56, 57, 26, 58], DM self-interactions [59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 15, 17], as well as warm DM models [69, 70, 71, 72, 73, 74, 75, 76, 77, 78], and axion DM models [79, 80, 81, 82, 83].
@@ 2 Dark matter-neutrino interactions
In this work we assume a scattering interaction between dark matter and standard model neutrinos, as described in Refs. [42, 43, 44]. A coupling between neutrinos and dark matter has been suggested as a possible source of the non-zero neutrino masses, and such couplings are exceedingly difficult to probe in laboratory experiments because of the weak coupling between neutrinos and other standard model particles [115, 116, 117]. We use the approximation of massless neutrinos for the sake of simplicity in the EFTofLSS calculations, but this is unlikely to have a large impact on the final constraints, as the inclusion of neutrino masses has been shown to only have minor effects on the bounds of the interaction strength [45, 48].
@@
$u_{\nu\chi}\equiv\frac{\sigma_{\nu\chi}}{\sigma_{\mathrm{Th}}}\left(\frac{m_{\chi}}{100\,\mathrm{GeV}}\right)^{-1},$ (2.1)
where $\sigma_{\mathrm{Th}}\approx 6.65\times 10^{-29}\,\mathrm{m}^{2}$ is the Thomson scattering cross-section, $m_{\chi}$ is the dark matter particle mass, and $\sigma_{\nu\chi}$ is the elastic scattering cross section between interacting DM particles and neutrinos. As it describes a scattering probability, it will always be positive or zero. The mass factor is included in the definition of the interaction strength to avoid a parameter degeneracy between the scattering cross section and the DM particle mass, as the interaction rate is proportional to the scattering cross section and the DM particle number density.
The updated Euler equations for dark matter and neutrinos can be derived from the matrix element of the coupling following the approach of ref. [114], and are then given as [27, 40, 41, 42, 43, 44]
$\dot{\theta}_{\nu}$ $=k^{2}\psi+k^{2}\left(\frac{1}{4}\delta_{\nu}-\sigma_{\nu}\right)-\dot{\mu}\left(\theta_{\nu}-\theta_{\chi}\right),$ (2.2a)
$\dot{\theta}_{\chi}$ $=k^{2}\psi-\mathcal{H}\theta_{\chi}-\frac{4\rho_{\nu}}{3\rho_{\chi}}\dot{\mu}\left(\theta_{\chi}-\theta_{\nu}\right),$ (2.2b)
where $\theta_{\nu}$ and $\theta_{\chi}$ are the neutrino and interacting DM velocity divergences, $k$ is the comoving wave number, $\psi$ is the Newtonian gauge gravitational potential, $\mathcal{H}$ is the conformal Hubble parameter, $\rho_{\nu}$ and $\rho_{\chi}$ are the neutrino and interacting DM energy densities, and $\delta_{\nu}$ and $\sigma_{\nu}$ are the neutrino density perturbation and anisotropic stress, respectively. The interaction rate $\dot{\mu}$ is given through,
$\dot{\mu}\equiv a\sigma_{\nu\chi}n_{\chi}=\frac{\sigma_{\mathrm{Th}}}{100\,\mathrm{GeV}}au_{\nu\chi}\rho_{\chi},$ (2.3)
where $n_{\chi}$ is the interacting DM particle number density.
We restrict our analysis to the simplest form of interaction, using a temperature independent cross section. In terms of the underlying particle physics model, this matches a subset of the simplified models presented in ref. [117], namely models with a scalar DM and a Fermionic mediator, or vector DM with a Dirac mediator, where the DM and mediator particles are mass degenerate. There are no specific assumptions on the mass of the DM particle other than it has to be heavy enough that it can be approximated to be at rest for the interaction, i.e. it is non-relativistic and has a mass much larger than the neutrino energy.
Figure 1: Linear and EFTofLSS matter power spectrum at redshift $z=0$ for two IDM scenarios, with 50% and 90% IDM, respectively, as well as standard CDM for reference. For the nonlinear power spectrum we have set counterterms to zero. It is evident that the scenario with the larger IDM fraction is more suppressed, but for both models the suppression starts at the same scale, as they have the same interaction strength $u_{\nu\chi}$.
@@ 3 Effective Field Theory of Large Scale Structure
For the definition of EFTofLSS parameters, we rely on the CLASS-OneLoop parameterization, which is based on the convention in [121] and detailed in [109]. When modelling the one-loop galaxy power spectrum in real space at a given redshift, we have four bias parameters $(b_{1},b_{2},b_{{\cal G}_{2}},b_{\Gamma_{3}})$, one counter-term $c_{0}$, and two stochastic parameters $(s_{0},s_{1})$ For the one-loop galaxy power spectrum wedges in redshift space, we have the same four bias parameters, four counter-terms $(c_{0},c_{1},c_{2},c_{3})$, and four stochastic terms $(s_{0},s_{1},s_{2},s_{3})$. [^3] Ref. [109] showed that $s_{1}$ is mostly degenerate with other parameters. Thus, in the following, we can limit ourselves to varying the three parameters $(s_{0},s_{2},s_{3})$. Note that $s_{3}$ is often refereed to as $s_{4}$ in other works.
@@
with flat uninformative priors on $(\tilde{b}_{1},\tilde{b}_{2},\tilde{b}_{{\cal G}2},\tilde{b}_{\Gamma_{3}})$. The fiducial values of $b_{{\cal G}_{2}},b_{\Gamma_{3}}$ are chosen according to the co-evolution model while $b_{2}$ follows a polynomial fit; this is the same approach as used in [109]. This co-evolution picture assumes that there were initially no biases associated with the tidal field. This implies that the late-time tidal biases ($b_{{\cal G}_{2}}$,$b_{\Gamma_{3}}$) are fully determined by local-in-matter biases ($b_{1}$,$b_{2}$) through the late-time gravitational evolution [130]. Since the interaction affects relatively small scales (comoving scales smaller than the sound horizon of the coupled fluid at DM-neutrino kinetic decoupling [^4] The exact decoupling time, and thus suppression scale, is set by the interaction strength, with a weaker interaction decoupling earlier and thus having a suppression scale at larger $k$. See Fig. 1 for an example.) and decouples long before the onset of non-linear structure formation, it neither modifies the dynamics on large scales, nor induces a scale-dependency within the growth, thus we might assume the co-evolution to hold similarly to $\Lambda$CDM in our case. For the seven counter-terms and stochastic parameters, we compare three assumptions:
@@ 4.2 Comparison with intermediate and pessimistic cases
Overall, the impact of loosening our assumptions on stochastic terms and counterterms indicates that there is a significant degeneracy between the interaction parameter $u_{\nu\chi}$ and the stochastic contributions or counterterms. Our results in the intermediate and the pessimistic cases indicate that in the absence of well-motivated informative priors on these nuisance parameters, the limits on dark matter-neutrino interactions from galaxy clustering power spectrum using the EFTofLSS model will likely not to be competitive with those from the CMB. If the assumptions on the redshift evolution of the four bias parameters were loosened, this could lead to similar degradation in constraining power. This conclusion is expected to be extendable to other models with small-scale suppression of linear matter power spectrum. However, the constraining power observed in the optimistic case suggests that improved constraints may be achievable in the future provided that a physically-motivated model for the redshift dependence of the EFTofLSS parameters can be established.
@@ Acknowledgments
The authors would like to thank Felix Kahlhöfer for useful discussions. The authors gratefully acknowledge the computing time provided to them on the high-performance computer Lichtenberg at the NHR Center NHR4CES@TUDa, funded by the German Federal Ministry of Education and Research (BMBF) and the Hessian Ministry of Science and Research, Art and Culture (HMWK). JL and MM acknowledge support from the DFG grant LE 3742/8-1. Some tests of the likelihood were performed with computing resources granted by RWTH Aachen University under projects thes1654 and rwth1389. SC acknowledges support by the Science and Technology Facilities Council (grant number ST/W001225/1). The authors gratefully acknowledge the computing time provided to them at the NHR Center NHR4CES at RWTH Aachen University (project number p0021792). This is funded by the Federal Ministry of Education and Research, and the state governments participating on the basis of the resolutions of the GWK for national high performance computing at universities (www.nhr-verein.de/unsere-partner).